OpenStudy (anonymous):

which function at the bottom is the inverse of f(x) = x2 − 36? x squared over 36 ±6square root of x 1 over quantity x squared minus 36 ±square root of the quantity x plus 36 need help

OpenStudy (luigi0210):

To solve for an inverse switch y and x and then solve for y

OpenStudy (anonymous):

I've tried that and keep coming up with the wrong answer

OpenStudy (luigi0210):

Well show us your work and process, we will try to see where you messed up

OpenStudy (anonymous):

1.)let f(x)=y 2.) switch x and y 3.)solve for y( rearrange the equation in terms of y)

OpenStudy (anonymous):

f(x)=x2-36 y=x2-36 x=y2-36 ??

OpenStudy (zpupster):

ok whats next--isolate the y

OpenStudy (anonymous):

how?

OpenStudy (zpupster):

add 36 to both sides

OpenStudy (zpupster):

what did you get??

OpenStudy (anonymous):

x+36=y2?

OpenStudy (zpupster):

ok how would get y by itself

OpenStudy (anonymous):

that's what i'm trying to remember lol

OpenStudy (zpupster):

sqrt both sides

OpenStudy (anonymous):

ok thanks

OpenStudy (zpupster):

good job!!!!!

OpenStudy (tkhunny):

An Inverse Function or just any Inverse Relation? I do wish these questions would be more specific.